3g C completely reacts with 5g O2 what is the %age concentration  of CO in the product?? 

Grade:11

1 Answers

Arun
5 years ago
 
Solution:
The reaction for the same would be given by:
3C+2O2\rightarrow 2CO+CO2
Now the number of moles of C are equal to:
\frac{3}{12}=\frac{1}{4}
The number of moles of Oxygen are:
\frac{5}{16}
Thus the limiting reagent is C.
And so the mole fraction of CO2 would be equal to:
\frac{\frac{1}{3}}{\frac{1}{3}+\frac{2}{3}}=\frac{1}{3}